# Mixed precision iterative refinement for least squares with linear equality constraints and generalized least squares problems

### v2

@@ 1 Introduction
      The solution $x$ is unique for any $b\in\mathbb{C}^{m}$. This is satisfied if and only if $\mathcal{N}(A)\cap\mathcal{N}(B)=\{0\}$, which is equivalent to
@@
    The LSE problem is particularly important in applied regression analysis, one of the universal methods of mathematical modelling. Note that normal regression analysis takes the form of $y=X\beta+e$, which develops into a least squares problem of $\min_{\beta}\lVert X\beta-y\rVert_{2}$. If there exist certain equality constraints on $\beta$, e.g., from an exact observation or a physical law, then the standard least squares problem becomes an LSE problem (1.1). Therefore it is frequently encountered in applications such as constrained surface fitting, constrained optimization, geodetic least squares adjustment, signal processing, and other fields of interest; see [6, 16] for more details.
@@
    The GLS problem: Given matrices $W\in\mathbb{C}^{n\times m}$ and $V\in\mathbb{C}^{n\times p}$ with $m\leq n\leq m+p$ and vectors $d\in\mathbb{C}^{n}$, the GLS problem is described as
@@
    $\min_{x\in\mathbb{C}^{m},\,y\in\mathbb{C}^{p}}\Big{\lVert}[0,I_{p}]\begin{bmatrix}x\\ y\end{bmatrix}\Big{\rVert}_{2}\quad\text{s.t.}\quad[W,V]\begin{bmatrix}x\\ y\end{bmatrix}=d.$ (1.4)
@@
    $\operatorname{rank}\bigl{(}[W,V]\bigr{)}=n,\qquad\operatorname{rank}(W)=m,$ (1.5)
@@
    The GLS problem is crucial in economics and engineering as it is directly associated with the generalized linear regression model (GLM). GLM is a generalization of the standard linear regression model. Given a matrix $W\in\mathbb{C}^{n\times m}$ with full column rank, the GLM is represented as
@@
    where $e$ is a random vector with mean 00 and covariance matrix $\sigma^{2}Y$. If $\operatorname{rank}(Y)=p$, then $Y$ can be factorized to $Y=VV^{\top}$, where $V\in\mathbb{C}^{n\times p}$ has linearly independent columns. Thus (1.6) can be rewritten as
@@
    where $y$ is a random vector with mean 00 and covariance matrix $\sigma^{2}I$. The GLS problem (1.3) is raised to minimize $\sigma^{2}$. We refer readers to [30, 31] for more details.
@@
Mixed precision algorithms have attracted great attention in recent years partly due to the evolution of machine learning techniques and hardware infrastructure. The industry are not satisfied with performing all floating-point operations in a fixed precision (a.k.a. working precision), as lower precision arithmetic conducts faster floating-point operations and requires less data transfer. However, exploiting lower precision arithmetic may reduce the accuracy of the final solution. A clever use of lower precision arithmetic offers the possibility of greatly accelerating the computation, while still delivering an accurate solution in the working precision at the end. This paves the way for mixed precision algorithms in various applications; see the surveys [1, 21].
@@
Existing works on mixed precision algorithms in numerical algebra have already covered a number of topics. Those include algorithms for solving linear systems [9, 10, 24] as well as algorithms for solving least squares problems [11]. A brief investigation into half precision (FP16) arithmetic is discussed in [17, 18]. Topics related to eigenvalues and singular values are more complicated. Still, there exist several studies on this matter, including the symmetric eigensolvers [15, 23, 26, 27, 28] and nonsymmetric eigensolvers [7].
@@
Let us focus more closely on linear systems and least squares problems. The modern mixed precision approach for solving linear systems has been adequately presented by Carson and Higham in [10], which calculates the most expensive part—LU factorization—in lower precision and uses the classical iterative refinement algorithm to refine the solution to the working precision. For ill-conditioned matrices, however, this approach may fail to converge. Therefore [10] applies a GMRES solver, capable of utilizing the lower precision solution as the initial guess and lower precision LU factors as preconditioner, in such scenarios to further expand the feasibility of the algorithm. Since least squares problems can be transformed to certain linear systems, mixed precision techniques for linear systems carry over to least squares problems. Carson, Higham, and Pranesh proposed a mixed precision GMRES-based iterative refinement algorithm by performing a lower precision QR factorization to compute the initial guess and then reusing the QR factors as preconditioner to solve least squares problems through augmented systems [11, 22]. Mixed precision algorithms on the weighted least squares and total least squares problems have been studied by Carson and Oktay in [12] and [29], respectively.
@@
In this paper, we propose mixed precision classical and GMRES-based iterative refinement algorithms for LSE and GLS problems by utilizing the GRQ or GQR factorization. We analyse the four-precision iterative refinement algorithm and prove that the mixed precision algorithm is able to refine the lower precision solutions to the working precision level. We also show that when employing a higher precision in the computation of residuals, the iterative refinement method is able to improve the accuracy of the solution. Our numerical experiments further indicate that for well-conditioned problems, these mixed precision algorithms offer significant speedup compared to the fixed precision algorithm without sacrificing accuracy.
@@
[…]
@@ 2.1 Algorithms for the LSE problem
For the LSE problem (1.1), we first present the null-space approach via the GRQ factorization proposed by [2, 6] as follows:
@@
    where $R\in\mathbb{C}^{p\times p}$, $T\in\mathbb{C}^{m\times n}$ are upper triangular matrices, and $Z\in\mathbb{C}^{m\times m}$, $Q\in\mathbb{C}^{n\times n}$ are unitary matrices.
@@ 2.2 Algorithms for the GLS problem
As shown in Section 2.1, the null-space approach utilizes the GRQ factorization to solve the LSE problem. Similarly, Paige [30, 31] proposed an algorithm to address the GLS problem through the GQR factorization as follows:
@@
    Compute $x$ by solving the triangular system $Rx=c_{1}-T_{12}s_{2}$.
@@ 3 Mixed precision algorithms
Let $\bm{u}$ be the unit roundoff at the working precision. Similarly to [11], we employ four precisions $1\gg\,\bm{u}_{f}\geq\bm{u}_{s}\geq\bm{u}\geq\bm{u}_{r}$ in our analysis, where $\bm{u}$ is the working precision, $\bm{u}_{f}$, $\bm{u}_{r}$, and $\bm{u}_{s}$ stand for the unit roundoff used in the matrix factorization, the computation of residuals, and the correction equation solver, respectively. The general framework of mixed precision LSE/GLS algorithms is as follows:
@@
For LSE and GLS problems, as discussed in Section 2, we shall use the null-space approach and Paige’s algorithm, respectively, to perform the first step of the process at precision $\bm{u}_{f}$ so as to obtain the initial guess and the GRQ/GQR factorization. In this section, we further develop the four-precision classical iterative refinement method.
@@ 3.1 Classical iterative refinement for the LSE problem
where $R_{B}=[0,R]$; $Q\in\mathbb{C}^{n\times n}$ and $Z\in\mathbb{C}^{m\times m}$ are unitary matrices, $R\in\mathbb{C}^{p\times p}$ and $T\in\mathbb{C}^{m\times n}$ are upper triangular matrices, we compute the residual $r_{0}=b-Ax_{0}$ and $v_{0}$ by solving the triangular system $R^{*}v_{0}=(QA^{*}r_{0})(n-p+1:n)$ at precision $\bm{u}_{f}$, respectively.
@@
0:  A GRQ factorization of $(B,A)$ with upper triangular matrices $R\in\mathbb{C}^{p\times p}$, $T\in\mathbb{C}^{m\times n}$, and unitary matrices $Z\in\mathbb{C}^{m\times m}$, $Q\in\mathbb{C}^{n\times n}$; vectors $f_{1}\in\mathbb{C}^{m}$, $f_{2}\in\mathbb{C}^{p}$, and $f_{3}\in\mathbb{C}^{n}$.
@@
0:  A vector $x\in\mathbb{C}^{n}$ storing at precision $\bm{u}$ approximating the solution of the LSE problem (1.1).
@@
7:     if the stopping criteria (3.3) is satisfied then
@@
And in each refinement step, we have
@@ 3.2 Classical iterative refinement for the GLS problem
This process for the GLS problem concludes when the following stopping criteria are satisfied, with $\texttt{tol}=\mathcal{O}(\bm{u})$.
@@
Note that (3.8) can be solved through the same order as in Section 3.1. This procedure is formally outlined in Algorithm 3. Incorporating Algorithm 3 with the mixed precision framework discussed in the beginning of the section, we obtain the mixed precision GLS algorithm as in Algorithm 4.
@@
0:  A GQR factorization of $(W,V)$ with upper triangular matrices $R\in\mathbb{C}^{m\times m}$, $T\in\mathbb{C}^{n\times p}$, and unitary matrices $Z\in\mathbb{C}^{p\times p}$, $Q\in\mathbb{C}^{n\times n}$; vectors $f_{1}\in\mathbb{C}^{p}$, $f_{2}\in\mathbb{C}^{n}$, and $f_{3}\in\mathbb{C}^{m}$.
@@
0:  Vectors $x\in\mathbb{C}^{m}$ and $y\in\mathbb{C}^{p}$ storing at precision $\bm{u}$ approximating the solution of the GLS problem (1.3).
@@
6:     if the stopping criteria (3.7) is satisfied then
@@
for each refinement step, where $t_{f}$, $t_{w}$, $t_{r}$, and $t_{s}$ represent the total number of floating-point operations at precision $\bm{u}_{f}$, $\bm{u}$, $\bm{u}_{r}$, and $\bm{u}_{s}$, respectively. If $p\geq n\geq m$, we obtain
@@ 4 Accuracy analysis
We first introduce some notation regarding rounding error analysis. For simplicity, we use ^$\,\hat{\cdot}=\operatorname{f{}l}(\cdot)$ to represent the computed values interchangeably without further clarification. Let $\gamma_{n}=n\bm{u}/(1-n\bm{u})$ and $\tilde{\gamma}_{n}=cn\bm{u}/(1-cn\bm{u})$, where $c$ is a small constant. Here we assume $n\bm{u}\ll 1$. The superscripts $f$, $s$, and $r$ stand for the procedure computed at precision $\bm{u}_{f}$, $\bm{u}_{s}$, and $\bm{u}_{r}$, respectively.
We start by providing some assumptions and notations used in our analysis in Section 4.1. Then we shall present our analysis in several parts. In Section 4.2, we introduce a lemma on the accuracy of the GRQ/GQR factorization. Then in each subsection we conduct a rounding error analysis of our mixed precision algorithm. We summarize our results in Theorems 4.7 and 4.9 and derive the limiting accuracy in Theorems 4.8 and 4.10, respectively. We then present several examples to illustrate our findings. The analysis is partly based on the works of Björck in [5].
@@ 4.1 Basic assumptions
Assume that $A\in\mathbb{C}^{m\times n}$, $x\in\mathbb{C}^{n}$, and $y\in\mathbb{C}^{m}$. For the matrix–vector multiplication, i.e., $\mathtt{xGEMV}$ in BLAS/LAPACK [3], the computed result $\hat{y}=\operatorname{f{}l}(Ax)$ satisfies
@@
where $\epsilon_{\mathtt{gemv}}(n)$ is a function of $\bm{u}$ and $n$.
@@
Assume that $A\in\mathbb{C}^{m\times n}$. For the QR factorization $A=QR$, i.e., $\mathtt{xGEQRF}$ in LAPACK [3], with unitary $Q\in\mathbb{C}^{m\times m}$ and triangular $R\in\mathbb{C}^{m\times n}$, there exists a unitary matrix $Q_{0}\in\mathbb{C}^{m\times m}$ such that
@@
where $\hat{R}$ is the computed triangular factor, $A=[a_{1},\dotsc,a_{n}]$, and $\epsilon_{\mathtt{qr}}(m,n)$ denote a function of $\bm{u}$, $m$ and $n$.
@@
Assume that $T\in\mathbb{C}^{n\times n}$ is a triangular matrix and $b\in\mathbb{C}^{n}$. For the triangular system solver $Tx=b$, i.e., $\mathtt{xTRSV}$ in BLAS/LAPACK [3], the computed result $\hat{x}$ satisfies
@@
where $\epsilon_{\mathtt{trsv}}(n)$ is a function of $\bm{u}$ and $n$.
@@
Assume that $Q\in\mathbb{C}^{m\times m}$ is a unitary matrix and $y$, $z\in\mathbb{C}^{m}$. For applying $Q$ to $z$, i.e., $\mathtt{xORMQR}/{}\mathtt{xUNMQR}$ in LAPACK [3], the computed result $\hat{y}=\operatorname{f{}l}(Qz)$ satisfies
$\hat{y}=Q(z+\Delta z),\qquad$ $\lvert\Delta z\rvert\leq\epsilon_{\mathtt{unmqr}}(m)\lvert z\rvert,$
$\hat{y}=Qz+\Delta y,\qquad$ $\lvert\Delta y\rvert\leq\epsilon_{\mathtt{unmqr}}(m)\lvert\hat{y}\rvert,$
where $\epsilon_{\mathtt{unmqr}}(m)$ denote a function of $\bm{u}$ and $m$.
In Table 1, we present several established results related to these $\epsilon$. Nonetheless, stricter bounds are available, prompting the use of the more general notation $\epsilon_{*}(n)$. To simplify, we sometimes omit $n$ in $\epsilon_{*}(n)$ when the dimension size is unambiguous.
Table 1: Notation of rounding error analysis. Here $A\in\mathbb{C}^{m\times n}$; $Q$, $Q_{0}\in\mathbb{C}^{m\times m}$ are unitary matrices; $T\in\mathbb{C}^{n\times n}$ and $R\in\mathbb{C}^{m\times n}$ are triangular matrices; $b$, $x\in\mathbb{C}^{n}$ and $y$, $z\in\mathbb{C}^{m}$ are vectors.
@@
$\epsilon_{\mathtt{gemv}}(n)$ xGEMV $y=Ax$ $\gamma_{n}$ [20, Sect. 3.5]
$\epsilon_{\mathtt{qr}}(m,n)$ xGEQRF $A=QR$ $\tilde{\gamma}_{mn}$ [20, Sect. 19.3]
$\epsilon_{\mathtt{trsv}}(n)$ xTRSV $Tx=b$ $\gamma_{n}$ [20, Sect. 8.1]
$\epsilon_{\mathtt{unmqr}}(m)$ xUNMQR $y=Qz$ $\tilde{\gamma}_{m}$ [5] and [20, Sect. 19.5]
Employing pre-defined notations from Table 1 instead of existing error bounds like $\gamma_{n}$ is beneficial to the discussion of rounding errors. On one hand, it helps us to trace back and locate where the dominant rounding error terms come from and provides better insight for the mixed precision algorithm. On the other hand, bounds like $\gamma_{n}$ are often an overestimate. It may fail to provide useful information if the matrix dimension $n$ is large and the working precision is too low. Therefore it is recommended to use such pre-defined notations in future rounding error analysis.

### v3

@@ 1 Introduction
      The solution $x$ is unique for any $b\in\mathbb{C}^{m}$. This is satisfied if and only if $\mathcal{N}(A)\cap\mathcal{N}(B)=\{0\}$, where $\mathcal{N}(\cdot)$ represents the nullspace of a matrix, i.e., $\mathcal{N}(A)=\{x\colon Ax=0\}$. This is equivalent to
@@
    The LSE problem is particularly important in applied regression analysis, one of the universal methods of mathematical modelling. Note that normal regression analysis takes the form $b=Ax+e$, which transforms into the least squares problem $\min_{\beta}\lVert Ax-b\rVert_{2}$. If there exist certain equality constraints on $x$, e.g., from an exact observation or a physical law, then the standard least squares problem becomes an LSE problem (1.1). Therefore it is frequently encountered in applications such as constrained surface fitting, constrained optimization, geodetic least squares adjustment, signal processing, and other fields of interest. See [7, 18] for more details.
@@
    The GLS problem: Given matrices $W\in\mathbb{C}^{n\times m}$ and $V\in\mathbb{C}^{n\times p}$ with $m\leq n\leq m+p$ and vector $d\in\mathbb{C}^{n}$, the GLS problem is described as
@@
    $\min_{x\in\mathbb{C}^{m},\,y\in\mathbb{C}^{p}}\Big{\lVert}[0,I_{p}]\begin{bmatrix}x\\ y\end{bmatrix}\Big{\rVert}_{2}\qquad\text{s.t.}\qquad[W,V]\begin{bmatrix}x\\ y\end{bmatrix}=d.$ (1.4)
@@
    $\operatorname{rank}\big{(}[W,V]\big{)}=n,\qquad\operatorname{rank}(W)=m,$ (1.5)
@@
    The GLS problem is crucial in economics and engineering as it is directly associated with the generalized linear regression model (GLM). GLM is a generalization of the standard linear regression model. Given a matrix $W\in\mathbb{C}^{n\times m}$ with full column rank, the GLM is presented as
@@
    where $e$ is a random vector with mean 0 and covariance matrix $\sigma^{2}Y$. If $\operatorname{rank}(Y)=p$, then $Y$ can be factorized to $Y=VV^{*}$, where $V\in\mathbb{C}^{n\times p}$ has linearly independent columns. Thus (1.6) can be rewritten as
@@
    where $y$ is a random vector with mean 0 and covariance matrix $\sigma^{2}I$. The GLS problem (1.3) aims to minimize $\sigma^{2}$. We refer readers to [32, 33] for more details.
@@
Mixed precision algorithms have attracted great attention in recent years partly due to the evolution of machine learning techniques and hardware infrastructure. The industry is not satisfied with performing all floating-point operations in a fixed precision (a.k.a. working precision), as lower precision arithmetic conducts faster floating-point operations and requires less data transfer. However, exploiting lower precision arithmetic may reduce the accuracy of the final solution. A clever use of lower precision arithmetic offers the possibility of greatly accelerating the computation, while still delivering an accurate solution in the working precision at the end. This paves the way for mixed precision algorithms in various applications; see the surveys [1, 23].
@@
Existing works on mixed precision algorithms in numerical linear algebra have already covered a number of topics. Those include algorithms for solving linear systems [11, 12, 26] as well as algorithms for solving least squares problems [13]. A brief investigation into half precision (FP16) arithmetic is discussed in [19, 20]. Topics related to eigenvalues and singular values are more complicated. Still, there exist several studies on this matter, including the symmetric eigensolvers [17, 25, 28, 29, 30, 37] and nonsymmetric eigensolvers [8].
@@
Let us focus more closely on linear systems and least squares problems. The modern mixed precision approach for solving linear systems has been adequately presented by Carson and Higham in [12], which calculates the most expensive part—LU factorization—in lower precision and uses the classical iterative refinement algorithm to refine the solution to the working precision. For ill-conditioned matrices, however, this approach may fail to converge. Therefore [12] applies a GMRES solver, capable of utilizing the lower precision solution as the initial guess and lower precision LU factors as preconditioner, in such scenarios to further expand the feasibility of the algorithm. Since least squares problems can be transformed to certain linear systems, mixed precision techniques for linear systems carry over to least squares problems. Carson, Higham, and Pranesh proposed a mixed precision GMRES-based iterative refinement algorithm by performing a lower precision QR factorization to compute the initial guess and then reusing the QR factors as preconditioner to solve least squares problems through augmented systems [13, 24]. Mixed precision algorithms on weighted least squares [14] and total least squares problems [31] have been studied by Carson and Oktay.
@@
In this paper, we propose mixed precision classical and GMRES-based iterative refinement algorithms for LSE and GLS problems by utilizing the GRQ or GQR factorization. We analyse the four-precision iterative refinement algorithm and prove that the mixed precision algorithm is able to refine the lower precision solutions to the working precision level. We also show that when employing a higher precision in the computation of residuals, the iterative refinement method is able to improve the accuracy of the solution. Our numerical experiments further indicate that for well-conditioned problems, these mixed precision algorithms offer significant speedup compared to the fixed precision algorithms without sacrificing accuracy.
@@
We first introduce some notation used throughout the paper. We employ the MATLAB colon notation to present the submatrices and subvectors. Note that $\lVert\cdot\rVert_{2}$ indicates the $2$-norm, $\lVert\cdot\rVert_{\infty}$ indicates the infinity norm, and $\lVert\cdot\rVert_{\mathsf{F}}$ is utilized for the Frobenius norm. Let $\kappa_{2}(A)=\lVert A\rVert_{2}\cdot\lVert A^{\dagger}\rVert_{2}$ and $\kappa_{\infty}(A)=\lVert A\rVert_{\infty}\cdot\lVert A^{\dagger}\rVert_{\infty}$ denote the matrix condition numbers under different norms, where $A^{\dagger}$ represents the Moore–Penrose pseudoinverse of $A$. Additionally, $\lvert\cdot\rvert$ denotes the absolute value of each matrix element, and $\lvert A_{i,j}\rvert\leq\lvert B_{i,j}\rvert$ for all $i$ and $j$ is simplified as $\lvert A\rvert\leq\lvert B\rvert$. The notation $x\lesssim y$ signifies that $x$ is less than or equal to $y$ up to a multiplicative constant, i.e., $x\leq C\cdot y$ for some constant $C>0$.
@@ 2.1 Algorithms for the LSE problem
For the LSE problem (1.1), we first present the nullspace approach via the GRQ factorization proposed by [2, 7] as follows:
@@
    where $R\in\mathbb{C}^{p\times p}$ is an upper triangular matrix, $T\in\mathbb{C}^{m\times n}$ is an upper rectangular matrix, and $Z\in\mathbb{C}^{m\times m}$, $Q\in\mathbb{C}^{n\times n}$ are unitary matrices.
@@ 2.2 Algorithms for the GLS problem
As shown in Section 2.1, the nullspace approach utilizes the GRQ factorization to solve the LSE problem. Similarly, Paige [32, 33] proposed an algorithm to address the GLS problem through the GQR factorization as follows:
@@
    Compute $x\in\mathbb{C}^{m}$ by solving the triangular system $Rx=c_{1}-T_{12}s_{2}$.
@@ 3 Mixed precision algorithms
Let $\bm{u}$ be the unit roundoff at the working precision. Similarly to [13], we employ four precisions $1\gg\,\bm{u}_{f}\geq\bm{u}_{s}\geq\bm{u}\geq\bm{u}_{r}$ in our analysis, where $\bm{u}_{f}$, $\bm{u}_{r}$, and $\bm{u}_{s}$ stand for the unit roundoff used in the matrix factorization, the computation of residuals, and the correction equation solver, respectively. The general framework of mixed precision LSE/GLS algorithms is as follows:
@@
For LSE and GLS problems, as discussed in Section 2, we shall use the nullspace approach and Paige’s algorithm, respectively, to perform the first step of the process at precision $\bm{u}_{f}$ so as to obtain the initial guess and the GRQ/GQR factorization. In this section, we further develop the four-precision classical iterative refinement method.
@@ 3.1 Classical iterative refinement for the LSE problem
where $R_{B}=[0,R]$; $Q\in\mathbb{C}^{n\times n}$ and $Z\in\mathbb{C}^{m\times m}$ are unitary matrices, $R\in\mathbb{C}^{p\times p}$ is an upper triangular matrix, $T\in\mathbb{C}^{m\times n}$ is an upper rectangular matrix, we compute the residual $r_{0}=b-Ax_{0}$ and $v_{0}$ by solving the triangular system $R^{*}v_{0}=(QA^{*}r_{0})(n-p+1:n)$ at precision $\bm{u}_{f}$, respectively.
@@
0: A GRQ factorization of $(B,A)$ with the upper triangular matrix $R\in\mathbb{C}^{p\times p}$, the upper rectangular $T\in\mathbb{C}^{m\times n}$, and unitary matrices $Z\in\mathbb{C}^{m\times m}$, $Q\in\mathbb{C}^{n\times n}$; vectors $f_{1}\in\mathbb{C}^{m}$, $f_{2}\in\mathbb{C}^{p}$, and $f_{3}\in\mathbb{C}^{n}$.
@@
0: A vector $x\in\mathbb{C}^{n}$ stored at precision $\bm{u}$ approximating the solution of the LSE problem (1.1).
@@
7:  if the stopping criteria (3.3) are satisfied then
@@
And in each refinement step,
@@ 3.2 Classical iterative refinement for the GLS problem
This process for the GLS problem concludes when the following stopping criteria are satisfied, with $\mathtt{tol}=\mathcal{O}(\bm{u})$:
@@
Note that (3.8) can be solved through the same order as in Section 3.1. This procedure is formally outlined in Algorithm 3. Incorporating Algorithm 3 with the mixed precision framework discussed at the beginning of the section, we obtain the mixed precision GLS algorithm as in Algorithm 4.
@@
0: A GQR factorization of $(W,V)$ with the upper triangular matrix $R\in\mathbb{C}^{m\times m}$, the upper rectangular matrix $T\in\mathbb{C}^{n\times p}$, and unitary matrices $Z\in\mathbb{C}^{p\times p}$, $Q\in\mathbb{C}^{n\times n}$; vectors $f_{1}\in\mathbb{C}^{p}$, $f_{2}\in\mathbb{C}^{n}$, and $f_{3}\in\mathbb{C}^{m}$.
@@
0: Vectors $x\in\mathbb{C}^{m}$ and $y\in\mathbb{C}^{p}$ stored at precision $\bm{u}$ approximating the solution of the GLS problem (1.3).
@@
6:  if the stopping criteria (3.7) are satisfied then
@@
in each refinement step, where $t_{f}$, $t_{w}$, $t_{r}$, and $t_{s}$ represent the total number of floating-point operations at precision $\bm{u}_{f}$, $\bm{u}$, $\bm{u}_{r}$, and $\bm{u}_{s}$, respectively. If $p\geq n\geq m$, we obtain
@@ 4 Accuracy analysis
We first introduce some notation regarding rounding error analysis. For simplicity, we use $\hat{\cdot}=\operatorname{f{}l}(\cdot)$ to represent the computed values interchangeably without further clarification. Let $\gamma_{n}=n\bm{u}/(1-n\bm{u})$ and $\tilde{\gamma}_{n}=cn\bm{u}/(1-cn\bm{u})$, where $c$ is a small constant. Here we assume $n\bm{u}\ll 1$. The superscripts $f$, $s$, and $r$ stand for the procedure computed at precision $\bm{u}_{f}$, $\bm{u}_{s}$, and $\bm{u}_{r}$, respectively.
We start by providing some assumptions and notations used in our analysis in Section 4.1. Then we shall present our analysis in several parts. In Section 4.2, we introduce a lemma on the accuracy of the GRQ/GQR factorization. Then in each subsection we conduct a rounding error analysis of our mixed precision algorithm. We summarize our results in Theorems 4.7 and 4.9 and derive the limiting accuracy in Theorems 4.8 and 4.10, respectively. We then present several examples to illustrate our findings. The analysis is partly based on the work of Björck in [6].
@@ 4.1 Basic assumptions
Assume that $A\in\mathbb{C}^{m\times n}$, $x\in\mathbb{C}^{n}$, and $y\in\mathbb{C}^{m}$ so that $y=Ax$. For the matrix–vector multiplication, i.e., $\mathtt{xGEMV}$ in BLAS/LAPACK [3], the computed result $\hat{y}=\operatorname{f{}l}(Ax)$ satisfies
@@
where $\epsilon_{\mathtt{gemv}}(n)$ denotes a function of $\bm{u}$ and $n$.
@@
Assume that $A\in\mathbb{C}^{m\times n}$. For the QR factorization $A=QR$, i.e., $\mathtt{xGEQRF}$ in LAPACK [3], with unitary $Q\in\mathbb{C}^{m\times m}$ and upper rectangular $R\in\mathbb{C}^{m\times n}$, there exists a unitary matrix $Q_{0}\in\mathbb{C}^{m\times m}$ such that
@@
where $\hat{R}$ is the computed triangular factor, $A=[a_{1},\dotsc,a_{n}]$, and $\epsilon_{\mathtt{qr}}(m,n)$ denotes a function of $\bm{u}$, $m$ and $n$.
@@
Assume that $T\in\mathbb{C}^{n\times n}$ is a triangular matrix, and $b\in\mathbb{C}^{n}$. For the triangular system solver $Tx=b$, i.e., $\mathtt{xTRSV}$ in BLAS/LAPACK [3], the computed result $\hat{x}$ satisfies
@@
where $\epsilon_{\mathtt{trsv}}(n)$ denotes a function of $\bm{u}$ and $n$.
@@
Assume that $Q\in\mathbb{C}^{m\times m}$ is a unitary matrix, and $y$, $z\in\mathbb{C}^{m}$ so that $y=Qz$. For applying $Q$ to $z$, i.e., $\mathtt{xORMQR}/{}\mathtt{xUNMQR}$ in LAPACK [3], there exist a backward error $\Delta z$ and a forward error $\Delta y$ such that the computed result $\hat{y}=\operatorname{f{}l}(Qz)$ satisfies
$\hat{y}=Q(z+\Delta z),\qquad$ $\lVert\Delta z\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}(m)\lVert z\rVert_{2},$
$\hat{y}=Qz+\Delta y,\qquad$ $\lVert\Delta y\rVert_{2}\leq\epsilon_{\mathtt{unmqr}}(m)\lVert\hat{y}\rVert_{2},$
where $\epsilon_{\mathtt{unmqr}}(m)$ denotes a function of $\bm{u}$ and $m$.
In Table 1, we present several established results related to these $\epsilon_{*}(\cdot)$’s. Nonetheless, stricter bounds are available, prompting the use of the more general notation $\epsilon_{*}(\cdot)$.
Table 1: Notations of rounding error analysis. Here $A\in\mathbb{C}^{m\times n}$; $Q$, $Q_{0}\in\mathbb{C}^{m\times m}$ are unitary matrices; $T\in\mathbb{C}^{n\times n}$ is a triangular matrix; $R\in\mathbb{C}^{m\times n}$ is an upper rectangular matrix; $b$, $x\in\mathbb{C}^{n}$, and $y$, $z\in\mathbb{C}^{m}$ are vectors.
@@
$\epsilon_{\mathtt{gemv}}(n)$ xGEMV $y=Ax$ $\gamma_{n}$ [22, Section 3.5]
$\epsilon_{\mathtt{qr}}(m,n)$ xGEQRF $A=QR$ $\tilde{\gamma}_{mn}$ [22, Section 19.3]
$\epsilon_{\mathtt{trsv}}(n)$ xTRSV $Tx=b$ $\gamma_{n}$ [22, Section 8.1]
$\epsilon_{\mathtt{unmqr}}(m)$ xUNMQR $y=Qz$ $\tilde{\gamma}_{m^{2}}$ [22, Section 19.3]
Employing pre-defined notations from Table 1 instead of existing error bounds like $\gamma_{n}$ is beneficial to the discussion of rounding errors. On one hand, it helps us to trace back and locate where the dominant rounding error terms come from and provides better insight for mixed precision algorithms. On the other hand, bounds like $\gamma_{n}$ are often an overestimate. It may fail to provide useful information if the matrix dimension $n$ is large and the working precision is too low. Therefore it is recommended to use such pre-defined notations in future rounding error analysis.